# Towards an Information Theoretic Framework of Context-Based Offline Meta-Reinforcement Learning

### v1

@@ 4 Experiments
Our main experiments are organized to address three primary inquiries regarding the core advantages of UNICORN: (1) How does UNICORN perform on in-distribution tasks (bias-reduction)? (2) How well can UNICORN generalize to data collected by OOD behavior policies (robustness)? (3) Can UNICORN outperform consistently across datasets of different qualities (robustness)? The performance are measured by the average return across all test tasks.
Figure 3: The hyper-parameter plane of UNICORN. Empirically, we find that UNICORN works properly only when $(\alpha\beta-1)/(1-\alpha)\beta>0$, which further constrains the feasible region of UNICORN (shaded) to be bounded by $\alpha\beta\geq 1$, $\alpha\leq 1$ and $\beta\geq 1$. Intriguingly, pre-existing methods FOCAL (red) and CORRO/CSRO/GENTLE (blue) constitute the two boundaries of the feasible region, which can therefore be interpreted as special cases of UNICORN.
4.1 Environments and Datasets
We adopt MuJoCo (Todorov et al., 2012) and MetaWorld (Yu et al., 2020a) benchmarks to evaluate our method.
HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir are multi-task MuJoCo benchmarks where tasks differ in reward function. For HalfCheetah-Dir and Ant-Dir, the agent is rewarded with high velocity along the goal direction. For HalfCheetah-Vel, the agent receives an $L_{2}$ penalty with respect to the target velocity.
Hopper-Param, Walker-Param are multi-task MuJoCo benchmarks where tasks differ in transition dynamics. The physical parameters of body mass, inertia, damping, and friction vary for each task.
Reach is a multi-task MetaWorld benchmark where tasks differ in reward function. The agent controls the end-effector to reach a target position in 3D space.
Following the protocol of CORRO, we sample 20 training tasks and 20 testing tasks from the task distribution. We train SAC agent (Haarnoja et al., 2018) independently for each task and collect offline datasets of mixed quality using all logged behavior policies.
4.2 Baselines
We compare UNICORN with six competitive OMRL algorithms. These baselines are categorized as context-based, gradient-based or transformer-based methods.
FOCAL, CORRO, CSRO are context-based methods which can be seen as special cases or approximations of UNICORN, see details in Section 3.
@@ 4.2 Baselines
MACAW (Mitchell et al., 2021) is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
Prompt-DT (Xu et al., 2022) is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) (Chen et al., 2021; Janner et al., 2021) to solve OMRL as a conditional sequence modeling problem.
We perform our evaluation in a few-shot manner. For gradient-based methods, we first update the meta-policy with a batch of context, then evaluate the agent. For transformer-based methods, we utilize $\bm{z}$ inferred from a batch of context as the prompt to condition the auto-regressive rollout.
@@ 4.3 Few-Shot Generalization to In-Distribution Data
Algorithm HalfCheetah-Dir HalfCheetah-Vel Ant-Dir Hopper-Param Walker-Param Reach
IID OOD IID OOD IID OOD IID OOD IID OOD IID OOD
UNICORN 1312.59±31.72 1301.17±25.98 -21.88±0.65 -93.57±4.85 266.85±14.54 235.55±17.69 315.90±5.67 304.45±10.88 407.83±75.58 392.48±74.45 2774.67±241.04 2603.68±182.85
CSRO 1180.48±228.10 458.08±253.00 -27.98±0.90 -102.44±5.17 276.18±19.37 233.33±12.37 309.63±6.50 301.12±9.99 310.47±58.29 278.94±65.37 2719.50±234.55 2800.93±182.06
CORRO 704.20±449.46 244.95±145.83 -37.11±2.60 -112.29±1.62 147.93±12.85 120.52±12.48 283.04±7.89 271.48±12.79 277.07±37.91 213.43±48.33 2467.83±174.57 2321.87±327.19
FOCAL 1186.03±271.58 860.58±252.96 -22.22±0.88 -96.72±1.69 217.10±29.14 173.16±24.37 301.84±4.48 296.87±12.76 308.05±98.42 286.34±91.08 2423.67±255.95 2315.56±303.03
Supervised 961.62±355.84 781.51±429.12 -23.81±0.71 -104.31±0.65 237.52±38.81 202.19±37.56 306.54±9.99 293.65±7.99 256.52±60.39 209.64±27.72 2489.37±248.05 2283.26±205.36
MACAW 1154.83±10.11 449.56±6.54 -56.04±2.42 -188.44±0.91 26.29±2.80 0.18±0.31 218.52±6.01 205.39±2.47 141.25±8.83 129.63±5.09 2430.67±157.42 1728.01±78.98
Prompt-DT 1176.54±39.64 -25.10±8.74 -118.14±65.59 -248.64±21.38 1.00±0.38 0.04±0.11 234.19±5.35 202.51±4.96 184.73±8.57 155.78±16.64 2164.86±85.20 1895.51±110.64
Figure 4 illustrates the learning curves of UNICORN vs. all baselines. IID entries of Table 1 depicts the performance of in-distribution generalization where the context for each task is collected by the behavior policies trained on the same task. The results demonstrate that UNICORN almost unanimously achieves SOTA performance across all benchmarks. The observation that FOCAL, CORRO, and CSRO achieve asymptotic performance close to UNICORN is also expected, since they are special cases of UNICORN according to our theory. Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process. Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired. Moreover, the average training time for MACAW is almost three times longer than the context-based methods.
@@ 4.4 Few-Shot Generalization to Out-of-Distribution Behavior Policies
Denote by $\{\pi_{\beta}^{i,t}\}_{i,t}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration. For each test task $M^{j}$, we use every individual behavior policy in $\{\pi_{\beta}^{i,t}\}_{i,t}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1. The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 1. While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin. Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT, which suggests that context-based paradigm should be prioritized for research of generalizable OMRL. This also delineates the storyline of COMRL development from 2021 to date
@@ 4.5 Influence of Data Quality
Algorithm Random Medium Expert
IID OOD IID OOD IID OOD
UNICORN 81.31±17.80 62.20±6.15 219.86±23.04 242.74±10.40 278.82±10.46 261.83±13.49
CSRO 1.55±2.65 -0.20±0.52 165.65±9.89 198.15±17.35 251.77±39.16 201.57±45.29
CORRO 0.87±0.53 -0.05±0.06 8.16±5.02 -7.01±2.37 -3.90±9.59 -13.59±9.06
FOCAL 67.00±26.33 43.65±10.08 171.10±84.26 187.06±86.45 229.02±42.07 246.27±20.21
Supervised 64.97±5.57 46.90±11.86 149.21±50.21 109.85±80.09 249.11±33.47 214.55±59.56
MACAW 3.412±0.74 -0.27±0.25 27.69±2.17 1.34±0.75 87.86±42.69 0.64±1.49
Prompt-DT 0.54±0.29 0.02±0.09 2.14±4.36 0.33±0.57 77.94±14.67 1.24±2.18
To test whether the UNICORN framework can be applied to different data distributions, we collect three types of datasets whose size is equal to the mixed dataset used in Section 4.3 and 4.4: random, medium and expert, which are characterized by the quality of the behavior policies. Each dataset is collected by one SAC agent. For random datasets, the SAC agent is randomly initialized and untrained.
As shown in Table 2, UNICORN demonstrates SOTA performance on all types of datasets. Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed data distribution, UNICORN surpasses CSRO by a large margin under the narrow data distribution (medium and expert) and random data distribution.
@@ 5 Discussion
This section presents more empirical evidence on the broader applicability of the UNICORN framework.
@@ 5.1 Is UNICORN Model-Agnostic?
Since UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic. Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT. Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN}}$ and $\mathcal{L}_{\text{FOCAL}}$. We name these DT implementations as UNICORN-DT and FOCAL-DT respectively.
As shown in Table 3 and Figure 8, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option. Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer (Kaplan et al., 2020) and DT (Lee et al., 2022). We leave this verification to future work.
@@
Algorithm HalfCheetah-Dir Hopper-Param
IID OOD IID OOD
UNICORN 1312.59±31.72 1301.17±25.98 315.90±5.67 304.45±10.88
UNICORN-DT 1233.19±10.42 1185.76±42.71 303.66±3.51 290.88±3.68
FOCAL-DT 1208.58±33.38 652.14±35.86 293.25±4.17 283.86±4.81
Prompt-DT 1176.54±39.64 -25.10±8.74 234.19±5.35 202.51±4.96
@@ 5.2 Can UNICORN be Exploited for Model-Based Paradigms?
As alluded in Section 3.3, UNICORN can potentially enable a model-based paradigm by generating data using the decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$ and customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks. We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1). Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$. The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent. We adopt the conventional ensemble technique in model-based RL (Yu et al., 2020b; Zhang et al., 2023) to stabilize the training process.
As shown in Figure 5, UNICORN-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
@@
6 Conclusion
In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework based on the information bottleneck principle. Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift. Based on the proposed framework, we arrive at a novel algorithm called UNICORN which is demonstrated to be remarkably robust and broadly applicable through extensive experiments. We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully offline and offline-to-online COMRL paradigms. We leave these directions for future work.
@@ Appendix A Pseudo-Code
0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\omega}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, decoder $p_{\theta}$, hyper-parameter $\alpha$, $\beta$, task batch size $N_{tbatch}$
1:  while not done do
2:     Sample task batch $\{M^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{X^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
3:     for step in each iter do
4:        // Train the context encoder and decoder
5:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
6:        Obtain task embeddings $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
7:        Compute $\mathcal{L}_{\text{UNICORN}}=\mathcal{L}_{\text{recon}}+\frac{\alpha\beta-1}{(1-\alpha)\beta}\mathcal{L}_{\text{FOCAL}}$
8:        Update $\phi,\theta$ to minimize $\mathcal{L}_{\text{UNICORN}}$
9:        // Train the actor and critic
10:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
11:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
12:        Compute $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
13:        Update $\omega,\psi$ to minimize $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
14:     end for
15:  end while
@@
0:  Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{test}}$, testing tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{test}}$, learned policy $\pi_{\omega}$, context encoder $q_{\phi}$
1:  for each $M^{i}$ do
2:     Sample context $\bm{c}^{i}\sim X^{i}$
3:     Obtain task embedding $\bm{z}^{i}\sim q_{\phi}(\bm{z}|\bm{c}^{i})$
4:     Rollout policy $\pi_{\omega}(\bm{a}|\bm{s},\bm{z}^{i})$ for evaluation
5:  end for

### v2

@@ 3 Experiments
For brevity, we name our proposed supervised and self-supervised algorithms UNICORN-SUP and UNICORN-SS respectively. To evaluate their effectiveness, our main experiments are organized to address three primary inquiries regarding the core advantages of UNICORN: (1) How does UNICORN perform on in-distribution tasks? (2) How well can UNICORN generalize to data collected by OOD behavior policies? (3) Can UNICORN outperform consistently across datasets of different qualities? The performance is measured by the average return across 20 testing tasks randomly sampled for each environment.
3.1 Experimental Setup
We adopt MuJoCo [25] and MetaWorld [3] benchmark to evaluate our method, involving six robotic locomotion tasks which require adaptation across reward functions (HalfCheetah-Dir, HalfCheetah-Vel, Ant-Dir, Reach) or across dynamics (Hopper-Param, Walker-Param). We compare UNICORN with six competitive OMRL algorithms. These baselines are categorized as context-based, gradient-based and transformer-based methods:
Algorithm HalfCheetah-Dir HalfCheetah-Vel Ant-Dir Hopper-Param Walker-Param Reach IID OOD IID OOD IID OOD IID OOD IID OOD IID OOD UNICORN-SS 1307±26 1296±24 -22±1 -94±5 267±14 236±18 316±6 304±11 419±44 407±46 2775±241 2604±183 UNICORN-SUP 1296±20 1130±76 -25±3 -91±5 250±4 239±16 312±4 302±12 322±28 312±39 2681±111 2641±140 CSRO 1180±228 458±253 -28±1 -102±5 276±19 233±12 310±6 301±10 310±58 279±65 2720±235 2801±182 CORRO 704±450 245±146 -37±3 -112±2 148±13 120±12 283±8 272±13 277±38 213±48 2468±175 2322±327 FOCAL 1186±272 861±253 -22±1 -97±2 217±29 173±24 302±4 297±13 308±98 286±91 2424±256 2316±303 Supervised 962±356 782±429 -24±1 -104±1 238±39 202±38 306±10 294±8 256±60 210±28 2489±248 2283±205 MACAW 1155±10 450±6 -56±2 -188±1 26±3 0±0 218±6 205±2 141±9 130±5 2431±157 1728±79 Prompt-DT 1176±40 -25±9 -118±66 -249±21 1±0 0±0 234±5 202±5 185±9 156±17 2165±85 1896±111
FOCAL, CORRO, CSRO are context-based methods that can be seen as special cases or approximations of UNICORN, see details in Section 2.
@@ 3.1 Experimental Setup
MACAW [12] is a gradient-based method, using supervised advantage-weighted regression for both the inner and outer loop of meta-training.
Prompt-DT [42] is a transformer-based method, taking context as the prompt of a Decision Transformer (DT) [33, 34] to solve OMRL as a conditional sequence modeling problem.
We perform our evaluation in a few-shot manner. For gradient-based methods, we first update the meta-policy with a batch of testing context and then evaluate the agent. For transformer-based methods, we utilize the trajectory segment as the prompt to condition the auto-regressive rollout.
@@ 3.2 Few-Shot Generalization to In-Distribution Data
Denote by $\{\pi_{\beta}^{i,t}\}_{i=1:N_{M},t=0:N}$ the checkpoints of behavior policies for all tasks logged during data collection (i.e., training SAC agents), where $i$ refers to the task index and $t$ represents training iteration. In the in-distribution (IID) setting, for each test task $M^{j}$, we use behavior policies only from the same tasks $\{\pi_{\beta}^{j,t}\}_{j,t=0:N}$ to collect the context and the conditioned rollouts, as visualized in Figure 1. Figure 4 illustrates the learning curves of UNICORN vs. all baselines, which correspond to the IID entries of Table 2. The results demonstrate that UNICORN variants, especially UNICORN-SS, consistently achieve SOTA performance across all benchmarks. The observation that the asymptotic performance of CSRO is comparable to UNICORN is expected since it also optimizes a linear combination of the lower and upper bound of $I(\bm{Z};\bm{M})$ by Theorem 2.3, with a different implementation choice. Notably, UNICORN exhibits remarkable stability, especially on HalfCheetah-Dir and Walker-Param, where the other context-based methods suffer a significant performance decline during the later training process. Although the gradient-based MACAW and the transformer-based Prompt-DT show faster convergence, their asymptotic performance leaves much to be desired. Moreover, the average training time for MACAW is almost three times longer than the context-based methods.
@@ 3.3 Few-Shot Generalization to Out-of-Distribution Behavior Policies
To evaluate OOD generalization, for each test task $M^{j}$, we use all behavior policies in $\{\pi_{\beta}^{i,t}\}_{{i=1:N_{M},t=0:N}}$ to collect the OOD contexts and the conditioned rollouts, as visualized in Figure 1. The OOD testing performance is measured by averaging returns across all testing tasks, as shown in the OOD entries of Table 2. While all methods suffer notable decline when facing OOD context, UNICORN maintains the strongest performance by a large margin. Another observation is that the context-based methods are in general significantly more robust compared to the gradient-based MACAW and transformer-based Prompt-DT. This also delineates the storyline of COMRL development from 2021 to date
@@ 3.4 Influence of Data Quality
To test whether the UNICORN framework can be applied to different data distributions, we collect three types of datasets whose size is equal to the mixed dataset used in Section 3.2 and 3.3: random, medium and expert, which are characterized by the quality of the behavior policies.
As shown in Table 3, UNICORN demonstrates SOTA performance on all types of datasets. Notably, despite CSRO having better IID results than UNICORN on Ant-Dir under the mixed distribution, UNICORN surpasses CSRO by a large margin under the narrow distribution (medium and expert) and random distribution.
@@
Algorithm Random Medium Expert IID OOD IID OOD IID OOD UNICORN-SS 81±18 62±6 220±23 243±10 279±10 262±13 UNICORN-SUP 75±15 60±5 140±11 126±32 247±15 229±19 CSRO 2±3 0±1 166±10 198±17 252±39 202±45 CORRO 1±1 0±0 8±5 -7±2 -4±10 -14±9 FOCAL 67±26 44±10 171±84 187±86 229±42 246±20 Supervised 65±6 47±12 149±50 110±80 249±33 215±60 MACAW 3±1 0±0 28±2 1±1 88±43 1±1 Prompt-DT 1±0 0±0 2±4 0±1 78±15 1±2
@@ 4 Discussion
This section presents more empirical evidence on the applicability of the UNICORN framework.
@@ 4.1 Is UNICORN Model-Agnostic?
Algorithm HalfCheetah-Dir Hopper-Param IID OOD IID OOD UNICORN-SS 1307±26 1296±24 316±6 304±11 UNICORN-SS-DT 1233±10 1186±43 304±4 291±4 UNICORN-SUP-DT 1227±21 1065±57 308±6 297±2 FOCAL-DT 1209±33 652±36 293±4 284±5 Prompt-DT 1177±40 -25±9 234±5 203±5
As UNICORN tackles task representation learning from a general information theoretic perspective, it is in principle model-agnostic. Hence a straightforward idea is to transfer UNICORN to other model architectures like DT, which we envision to be a promising backbone for large-scale training of RL foundation models.
We employ a simple implementation by embedding the task representation vector $\bm{z}$ as the first token in sequence to prompt the learning and generation of DT. Instead of the unlearnable prompt used in the original Prompt-DT, we enforce task representation learning of the prompt by optimizing $\mathcal{L}_{\text{UNICORN-SS}}$, $\mathcal{L}_{\text{UNICORN-SUP}}$ and $\mathcal{L}_{\text{FOCAL}}$. We name these variants as UNICORN-SS-DT, UNICORN-SUP-DT and FOCAL-DT respectively.
As shown in Table 4 and Figure 7, applying UNICORN and FOCAL on DT results in significant improvement in both IID and OOD generalization, with UNICORN being the superior option. Despite the gap between our implementation of UNICORN-DT and its MLP counterpart in terms of asymptotic performance, we expect UNICORN-DT to extrapolate favorably in the regime of greater dataset size and model parameters due to the scaling law of transformer [43] and DT [44]. We leave this verification to future work.
@@ 4.2 Can UNICORN be Exploited for Model-Based Paradigms?
With decoder $p_{\theta}(\bm{x}_{t}|\bm{z},\bm{x}_{b})$, UNICORN-SS can potentially enable a model-based paradigm by generating data using customized latent $\bm{z}$, which can be especially useful for generalization to OOD tasks. We implement this idea on Ant-Dir by sorting the tasks according to the polar angle of their goal direction ($\theta\in[0,2\pi)$), and construct task-level OOD by taking 20 consecutive tasks in the middle for training and the rest for testing (Figure 1). Gaussian noise $\bm{\epsilon}$ is applied to the real task representation $\bm{z}$ and $(\bm{s},\bm{a},\bm{z}+\bm{\epsilon})$ is then fed to the decoder to generate $(\bm{s}^{\prime},r)$. The imaginary rollouts $\{(\bm{s}_{t},\bm{a}_{t},\bm{s}^{\prime}_{t},r_{t})\}_{t=1}^{n}$ are used during the training of RL agent. We adopt the conventional ensemble technique in model-based RL  [45, 46] to stabilize the training process.
As shown in Figure 5, UNICORN-SS-enabled model-based RL is the only method to achieve positive performance in this extremely challenging scenario where context shift is induced by OOD tasks.
5 Conclusion & Limitation
In this paper, we unify three major context-based offline meta-RL algorithms, FOCAL, CORRO and CSRO (potentially a lot more) into a single framework from an information theoretic representation learning perspective. Our theory offers valuable insight on how these methods are inherently connected and incrementally evolved in pursuit of better generalization against context shift. Based on the proposed framework, we instantiate two novel algorithms called UNICORN-SUP and UNICORN-SS which are demonstrated to be remarkably robust and broadly applicable through extensive experiments. We believe our study offers potential for new implementations, optimality bounds and algorithms for both fully-offline and offline-to-online COMRL paradigms. Mathematically, we expect our framework to be able to incorporate almost all existing COMRL methods that focus on task representation learning. Since according to Definition 2.1, as long as the method tries to solve COMRL by learning a sufficient statistic of w.r.t , it will eventually come down to optimizing an information-theoretic objective equivalent to $I(\bm{Z},\bm{M})$, or a lower/upper bound like the ones introduced by Theorem 2.3, up to some regularizations or constraints.
Limitation Since our framework assumes a decoupling of task representation learning and offline policy optimization, it does not directly elucidate how high-quality representations are indicative of higher downstream RL performance. Another limitation would be the scale of the experiments. Our offline datasets cover at most 40 tasks and $\sim 450k$ transitions, which might be the reason why UNICORN-SUP is inferior to UNICORN-SS ($n_{M}$ too small in Theorem 2.4). However, we expect our conclusion to extrapolate to larger task sets, dataset sizes and model parameters, for which the DT variants may demonstrate better scaling. Lastly, in principle, the information theoretic formalization in this work should be directly applicable to online RL. However, many of our key derivations rely on the static assumption of $\bm{M}$ and $\bm{X}$ (see Appendix B), which are apparently violated in the online scenario. Extending our framework to online RL is interesting and nontrivial, which we leave for future work.
Acknowledgments and Disclosure of Funding
The authors would like to thank the anonymous reviewers for their insightful comments and suggestions. The work described in this paper was supported partially by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N), by a grant from the Hong Kong Innovation and Technology Fund (Project Reference Number: MHP/086/21) and by the National Key Research and Development Program of China (No. 2020YFA0711402).
@@ Appendix A Pseudo-Code
1:Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{\textup{{train}}}}$, training tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{\textup{train}}}$, initialized learned policy $\pi_{\bm{\omega}}$, Q function $Q_{\bm{\psi}}$, context encoder $q_{\bm{\phi}}$, classifier/decoder $p_{\bm{\theta}}$, hyper-parameter $\alpha$, task batch size $N_{tb}$, learning rates $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$
2:while not done do
3:     Sample task batch $\{M^{j}\}_{j=1}^{N_{tb}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{X^{j}\}_{j=1}^{N_{tb}}\sim\bm{X}$
4:     for step in each iter do
5:         // Train the context encoder and decoder
6:         Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tb}}\sim\mathcal{R}$
7:         Obtain task embeddings $\{\bm{z}^{j}\sim q_{\bm{\phi}}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tb}}$
8:         Estimate $\mathcal{L}_{\text{UNICORN}}$ ($\mathcal{L}_{\text{UNICORN-SUP}}$ or $\mathcal{L}_{\text{UNICORN-SS}}$)
9:         $\bm{\phi}\leftarrow\bm{\phi}-\beta_{1}\nabla_{\bm{\phi}}\mathcal{L}_{\text{UNICORN}}$
10:         $\bm{\theta}\leftarrow\bm{\theta}-\beta_{2}\nabla_{\bm{\theta}}\mathcal{L}_{\text{UNICORN}}$
11:         // Train the actor and critic
12:         Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tb}}$
13:         Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tb}}\sim\mathcal{R}$
14:         Estimate $\mathcal{L}_{\text{actor}},\mathcal{L}_{\text{critic}}$
15:         $\bm{\omega}\leftarrow\bm{\omega}-\beta_{3}\nabla_{\bm{\omega}}\mathcal{L}_{\text{actor}}$
16:         $\bm{\psi}\leftarrow\bm{\psi}-\beta_{4}\nabla_{\bm{\psi}}\mathcal{L}_{\text{critic}}$
17:     end for
18:end while
@@
1:Offline Datasets $\bm{X}=\{X^{i}\}_{i=1}^{N_{\textup{test}}}$, testing tasks $\bm{M}=\{M^{i}\}_{i=1}^{N_{\textup{test}}}$, trained policy $\pi_{\bm{\omega}}$ and context encoder $q_{\bm{\phi}}$
2:for each $M^{i}$ do
3:     Sample context $\bm{c}^{i}\sim X^{i}$
4:     Obtain task embedding $\bm{z}^{i}\sim q_{\bm{\phi}}(\bm{z}|\bm{c}^{i})$
5:     Rollout policy $\pi_{\bm{\omega}}(\bm{a}|\bm{s},\bm{z}^{i})$ for evaluation
6:end for